An alloy of steel has a Young's modulus of 180GPa, elastic limit of 0.002 and ultimate tensile strength of 860MPa. An alloy of aluminium has a Young's modulus of 70GPa, elastic limit of 0.004 and ultimate tensile strength of 110MPa.
a) Sketch a stress-strain graph for both of these alloys, plotting them on the same graph
b) Label significant positions on the graphs
c) Compare and contrast the mechanical properties of these alloys
1
Expert's answer
2020-01-30T09:41:25-0500

a, b) Begin with significant points: elastic limit and Young's modulus. Take aluminum alloy as reference: its Young's modulus is 2.57 times smaller than that of steel , its elastic limit is 2 times longer, ultimate tensile strength of aluminum is 7.81 times smaller. Stress is strain times Young's modulus, therefore:


"\sigma_s=Y_s\epsilon_s=180\cdot10^9\cdot0.002=360\cdot10^6\text{ Pa},\\
\sigma_a=Y_a\epsilon_a=70\cdot10^9\cdot0.004=280\cdot10^6\text{ Pa}."

Take 860 MPa as 1:



c) Since the ultimate tensile strength of aluminum (green) is lower that its yield strength, it is nonsense, so perhaps there is a mistake in elastic limit or UTS. But as we see, aluminum can elastically elongate two times longer than steel. The stress steel experiences is higher.